OpenStudy (anonymous):

What is the angle of refraction of a light ray that enters the surface of a calm lake from air at an angle of 45o to the normal? The refractive index of water is 1.333. 32.04o 45.00o 59.99o 33.76o

OpenStudy (anonymous):

You use snells law in this problem \[n _{1} Sin \theta _{1} = n _{2}\sin \theta _{2}\]

OpenStudy (anonymous):

refractive index of air should be ridiculously close to 1. So you would have 1 * sin(45) = 1.333 * sin(theta) solve for theta